### v1

# PATCH: Learnable Tile-Level Hybrid Sparsity for LLMs

Abstract
Large language models (LLMs) deliver impressive performance but incur prohibitive memory and compute costs at deployment. Model pruning is an effective way to reduce these overheads, yet existing approaches face challenges: unstructured sparsity, where nonzeros can appear anywhere, preserves accuracy but yields irregular access patterns that prevent GPU acceleration, while semi-structured 2:4 sparsity is hardware-friendly but enforces a rigid 50% pattern that degrades model quality. To bridge this gap, we introduce PATCH, a hybrid sparsity framework that enables a continuous sparsity ratio between 0% and 50%. PATCH partitions weight matrices into tiles, assigning each tile to be either dense or 2:4 sparse via a learnable mask selection mechanism. This design provides fine-grained control over accuracy–acceleration tradeoffs and supports non-uniform sparsity across layers, leading to superior overall quality. Across models from 0.5B to 8B parameters, PATCH consistently narrows the gap to dense accuracy while delivering practical speedups. For instance, on LLaMA-2 7B with an A6000 GPU, PATCH achieves 1.18×–1.38× end-to-end speedup over dense baselines while improving accuracy by 0.37%–2.96% compared to the state-of-the-art 2:4 pruning method, MaskLLM.
1 Introduction
Recent advancements in large language models (LLMs) have revolutionized natural language processing, enabling breakthroughs in understanding and generating human language (Comanici et al., 2025; Meta, 2025). These models power diverse applications, such as conversational agents and automated content creation (Suzgun et al., 2022; Zhou et al., 2023). However, their extensive parameter counts—often in the billions—result in significant memory overhead and high inference costs (Guo et al., 2024; Ma et al., 2024). This computational burden has driven the need for efficient model compression techniques.
Two primary approaches to model compression are quantization and sparsity. Quantization reduces the precision of model parameters, compressing LLMs effectively while preserving performance (Ashkboos et al., 2024; Tseng et al., 2024; Zhang et al., 2024; Saha et al., 2024). In contrast, sparsity aims to lower memory and computational demands by setting many parameters to zero (Hassibi et al., 1993; LeCun et al., 1989). However, sparsity alone struggles to maintain model accuracy while delivering practical speedups, a limitation that current research seeks to overcome.
Unstructured sparsity, which permits non-zero elements to appear anywhere in the matrix, can match dense model accuracy due to its flexibility in sparsity allocation (Sun et al., 2023; Frantar & Alistarh, 2023; Agarwalla et al., 2024). However, its irregular memory access patterns hinder acceleration on modern hardware like GPUs (Xia et al., 2023; Fan et al., 2025). As a result, unstructured sparsity fails to deliver practical speedups, motivating the search for more hardware-friendly sparsity techniques.
Semi-structured sparsity patterns, such as the 2:4 pattern (Mishra et al., 2021) supported by NVIDIA and AMD GPUs, provide practical speedups in large-scale model inference. However, unlike unstructured sparsity, which offers greater flexibility, 2:4 enforces rigid rules by requiring at least two of every four consecutive elements to be zero. This rigidity often leads to significant accuracy loss when models are pruned using one-shot methods (Sun et al., 2023; Frantar & Alistarh, 2023; Ilin & Richtarik, 2025; Liu et al., 2025). MaskLLM (Fang et al., 2024) mitigates this issue by learning sparsity masks end-to-end, but pruned models still lag behind their dense counterparts in accuracy. Moreover, recent studies show that sparsity should be allocated non-uniformly (adaptively) across layers for optimal performance (Yin et al., 2025; Wang & Tu, 2020; Lee et al., 2021), whereas 2:4 sparsity enforces a fixed, uniform allocation. These limitations indicate that relying solely on 2:4 sparsity is insufficient, underscoring the need for hybrid approaches.
To address the challenges of LLM pruning, while providing accelerated inference, we propose Pruning with a Learnable Tile-level Configuration for Hybrid Sparsity (PATCH). PATCH learns a hybrid mask that partitions each weight matrix into hardware-friendly tiles, designating each tile as either dense (0% sparsity) or 2:4 sparse (50% sparsity). This adaptive mask allows the matrix to realize an effective global sparsity ratio anywhere between 0% and 50%, balancing accuracy in critical regions with hardware-friendly sparsity elsewhere. This design unites the hardware acceleration benefits of 2:4 sparsity with the flexibility of unstructured allocation, allowing sparsity to adapt to the varying importance of different layers. By jointly optimizing the sparsity within 2:4 tiles and the tile-level patterns during training, PATCH achieves higher accuracy than uniform sparsity across layers. Moreover, for resource-constrained settings, we offer a variant of PATCH that tunes only the dense tiles while freezing the initial 2:4 mask fixed. Importantly, PATCH is compatible with tile-level sparsity acceleration libraries and compilers such as STOICC (Rafii et al., 2025), making it the first hybrid sparsity method to demonstrate practical speedups. For example, on LLaMA-2 7B running on a consumer-grade A6000 GPU, PATCH achieves 1.18–1.38 end-to-end speedup over the dense baseline while improving accuracy by 0.37%–2.96% compared to the state-of-the-art 2:4 pruning method, MaskLLM.
2 Preliminaries
Differentiable Sampling.
Sampling from a categorical distribution is inherently non-differentiable, which poses challenges for gradient-based optimization. The Gumbel Softmax (Jang et al., 2016) addresses this by combining the Gumbel-Max reparameterization trick together with a softmax relaxation. The reparameterization expresses the sampling process by decoupling the deterministic log-probabilities from the stochastic perturbations introduced by Gumbel noise, which emulate random draws from the distribution. The subsequent softmax yields a differentiable approximation to categorical sampling:
where with . The resulting vector is a soft index vector whose entries represent the relaxed probability of selecting class .
Additionally, the temperature parameter controls the hardness of the sampled index. Lower values of yield a more peaked distribution, causing to converge to a one-hot vector as .
Learnable 2:4 Mask.
MaskLLM (Fang et al., 2024) formulates 2:4 mask selection as a learnable probabilistic process over the six possible patterns. The underlying weights remain fixed, while training shifts the categorical distribution to favor masks that preserve better pruning performance. The mask for each four consecutive elements can be parameterized with a vector . Scaling this vector to a weight matrix will result in as the mask search parameters. The resulting mask can be computed as in Equation 2, where denotes the 2:4 soft mask, obtained as a weighted average over the candidate masks, and is the matrix containing these six candidates as its rows.111We will refer to a mask value of as keeping the corresponding weight and a value of as pruning it.
A scaling factor is also introduced in Equation 1, where it multiplies the logits before adding the Gumbel noise , thereby controlling their relative influence. Small values let the noise dominate, encouraging exploration across candidate masks, while larger values amplify the logits and make the sampling more deterministic.
3 Patch
To overcome the rigidity of fixed 50% 2:4 sparsity, we introduce PATCH. PATCH learns a structured mask—optimized on top of frozen weights—that is partitioned into tiles, where each tile decides whether its corresponding weights remain dense or are pruned with a 2:4 pattern. This design preserves accuracy in sensitive regions while exploiting hardware-accelerated sparsity elsewhere. Unlike fixed 2:4 sparsity, which enforces the same pattern across all weights, PATCH adapts at the tile level by assigning dense tiles to critical regions and sparse tiles elsewhere.
Finding the optimal allocation of dense tiles (value 1) and sparse tiles (2:4 pattern) within a mask is a combinatorially difficult problem, as the number of possible configurations grows rapidly with the number of tiles across the LLM. By also modelling this problem as a probabilistic sampling process, and adjusting the probability of each tile (and the 2:4 patterns within sparse tiles), PATCH can efficiently explore the space of configurations and converge toward masks that balance accuracy and sparsity. The mask distributions are learned end-to-end by training the Gumbel–Softmax logits while keeping the model weights frozen. We address this challenge by formulating mask selection as two coupled subproblems: (1) selecting which tiles are dense or sparse, and (2) choosing the 2:4 sparsity pattern within sparse tiles.
Tile-based pruning of LLMs.
We associate each parameter matrix with a grid of tile-level distributions, each parameterized by a learnable logit. Collectively, these form , where each entry specifies the unnormalized score of keeping the corresponding tile fully dense. To create a two-class distribution (keep dense vs. prune), we concatenate a fixed zero to each logit, yielding . After applying Gumbel–Softmax, we broadcast the dense probabilities across their respective region (since the weighted average of the two outcomes reduces to ), so that all elements of a tile receive the same mask value. Formally,
This yields the tile-level mask in Equation 3, where is an all-ones matrix and denotes the Kronecker product.
Joint optimization with sparse mask.
To fully determine the effective sparsity pattern, the tile-level mask must be combined with the fine-grained 2:4 mask. Assuming that the 2:4 mask is generated using Equation 2, PATCH combines it with the tile mask as shown in Equation 4. The resulting soft mask interpolates between dense and sparse behavior: values of close to one make the tile predominantly dense, while values close to zero shift the tile toward the soft 2:4 mask pattern defined by . Thus, can be understood as a per-tile weighted average of the dense option and the 2:4 patterns, with determining the relative contribution of each. An overview of the process is provided in Figure 1.
Learning masks with targeted sparsity.
PATCH uses a novel regularization term to achieve a flexible 0%–50% sparsity ratio across the model by controlling the number of dense tiles. Unlike traditional regularization methods like weight decay, which produce non-deterministic sparsity ratios, our term penalizes deviations from the target sparsity, enabling precise control. This global sparsity approach prunes sensitive linear layers less aggressively while setting redundant weight elements to zero, offering greater flexibility than fixed per-layer sparsity. We directly compare global versus per-layer sparsity regularization in 5.
Training objective.
The overall training objective, as shown in Equation 5, of PATCH combines three components: the standard modeling loss, a sparsity regularization term that enforces the target density of the model , and a weight regularization term (as in MaskLLM) that promotes larger weight magnitudes and gradient propagation. Formally,
Following MaskLLM, we progressively decrease and increase during training so that the Gumbel-Softmax distribution converges to a clear one-hot choice of mask by the end of training.
Inference.
After training, the sign of each logit in determines the final mask. Since a zero logit is concatenated to represent the sparse class (Equation 3), positive values correspond to the dense option, while negative values correspond to the sparse option. The complete procedure is outlined in Algorithm 1.
Memory efficient PATCH.
To further reduce overhead, PATCH can be run in a memory-efficient manner by freezing the sparse mask parameters and optimizing only the tile-level decisions. This reduces the number of learnable parameters to . While this lighter formulation limits mask-selection flexibility and can reduce performance as seen in Table 5, it makes training feasible under strict memory constraints, such as fitting an 8B model on a single 80GB GPU. We denote this version of PATCH by PATCH and the joint optimization version of PATCH by PATCH.
4 Efficient deployment of PATCH
Executing PATCH requires handling hybrid sparse–dense tiles, a capability not supported by existing GPU libraries. Current tools either focus exclusively on dense computation (e.g., cuBLAS (NVIDIA Corporation, a), dense CUTLASS (Corporation, 2025), OpenAI Triton (Tillet et al., 2019)), or restrict support to fixed sparsity (e.g., cuSPARSELt (NVIDIA Corporation, b), sparse CUTLASS). STOICC (Rafii et al., 2025) lifts these limitations by extending Triton with hybrid tile-level sparsity, making it a suitable backend for accelerating PATCH.
Similar to Triton, STOICC employs an inspector that benchmarks candidate kernel configurations for each sparsity ratio, identifying the most hardware-efficient tile size for the target GPU. On NVIDIA A100 and A6000 GPUs, our experiments show that the optimal configurations are consistently drawn from or its subdivisions (e.g., , , ). In practice, this means that regardless of the sparsity ratio or the layer shape, the chosen granularity guarantees that STOICC’s autotuned tiles can be applied consistently. Unless otherwise specified, we adopt these hardware-friendly tile sizes in all PATCH experiments. Further implementation details are provided in Appendix A.
5 Experiments
Model, dataset and evaluation.
We evaluate PATCH across diverse transformer architectures, including the Qwen-2.5 (Qwen et al., 2025), Gemma 3 (Team et al., 2025), and LLaMA-2 (Touvron et al., 2023) and 3 (Grattafiori et al., 2024) model families, spanning 500M to 8B parameters. Following the dataset size and configurations in MaskLLM (Fang et al., 2024), masks are trained for 2000 steps with a batch size of 256 on sequences with a length of 4096 tokens from the SlimPajama dataset (Soboleva et al., 2023).
Following previous LLM compression work (Mozaffari et al., 2025a; Fang et al., 2024), we evaluate the models on eight zero-shot downstream tasks: PIQA (Bisk et al., 2020), ARC-Easy and ARC-Challenge (Clark et al., 2018), Winogrande (Sakaguchi et al., 2019), OpenBookQA (Mihaylov et al., 2018), RACE (Lai et al., 2017), HellaSwag (Zellers et al., 2019), and MMLU (Hendrycks et al., 2021) using the Language Model Evaluation Harness (Gao et al., 2024) framework. Additionally, similar to previous work (Mozaffari et al., 2025a; Frantar & Alistarh, 2023; Sun et al., 2023), we evaluate the models on a language modeling task using the WikiText2 (Merity et al., 2016) dataset with a sequence length of 4096, comparing against established baselines in the following sections.
Baselines.
To evaluate PATCH against established 2:4 sparsity pruning techniques, we compare it with the state-of-the-art learnable method MaskLLM (Fang et al., 2024), as well as one-shot methods including Wanda (Sun et al., 2023), SparseGPT (Frantar & Alistarh, 2023), Thanos (Ilin & Richtarik, 2025), ProxSparse (Liu et al., 2025) and magnitude pruning (Han et al., 2015). For one-shot pruning methods, following the default configurations in each paper, we prune the models over 128 samples from the C4 dataset.
The publicly available MaskLLM pruned checkpoints are limited to LLaMA-2 7B and LLaMA-3.1 8B models. To ensure a fair comparison across all models, we implemented MaskLLM in PyTorch and replicated its results for additional architectures presented in this study.
We faced a similar challenge with ProxSparse as well, where only the LLaMA-2-7B and LLaMA-3.1-8B checkpoints are publicly available. We have pruned other models with their official code base using their default hyperparameters for comparison.
Additional implementation details and hyperparameters used in our experiments are provided in Appendix D.
5.1 Model Quality Results
Joint sparse and dense tile optimization.
For smaller models like Qwen-2.5 0.5B, LLaMA-3.2 1B, and Gemma-3 1B, we apply the joint variant PATCH, which simultaneously optimizes dense tile locations and sparsity patterns within sparse tiles. This approach enables effective performance.
The average accuracy of the models across eight zero-shot downstream tasks and their perplexity on the WikiText2 dataset is reported in Table 1. The results demonstrate that PATCH provides a flexible tradeoff between sparsity ratio and model quality, narrowing the performance gap to dense models while ensuring hardware-friendly inference. A similar pattern holds for larger models using a memory-efficient variant, as explored next.
| Sparsity | Method | Pattern | Qwen-2.5 0.5B | LLaMA-3.2 1B | Gemma-3 1B | |||
| Acc (% ) | PPL () | Acc (% ) | PPL () | Acc (% ) | PPL () | |||
| 0% | Dense | - | 46.00 | 12.08 | 47.70 | 9.06 | 47.01 | 11.67 |
| 50% | Magnitude | 2:4 | 30.16 | 6734.97 | 29.66 | 563.44 | 31.66 | 5005.56 |
| Wanda | 2:4 | 32.97 | 72.48 | 31.61 | 78.18 | 34.16 | 69.41 | |
| SparseGPT | 2:4 | 34.81 | 36.59 | 35.55 | 32.73 | 35.58 | 44.59 | |
| Thanos | 2:4 | 31.31 | 37.32 | 35.71 | 33.03 | 35.09 | 62.63 | |
| ProxSparse | 2:4 | 32.05 | 111.05 | 33.55 | 49.33 | 36.63 | 90.50 | |
| MaskLLM | 2:4 | 39.33 | 15.22 | 41.04 | 12.93 | 41.84 | 12.82 | |
[…]
Memory-efficient tile selection.
For larger models such as LLaMA-2 7B and LLaMA-3.1 8B, we employ the memory-efficient variant PATCH, which freezes the fine-grained sparse weight structure while optimizing dense tile selections.
Table 2 summarizes the average accuracy of the models across eight downstream tasks in addition to their perplexity on the WikiText2 dataset for different sparsity ratios, illustrating that PATCH delivers a comparable flexible sparsity-quality tradeoff when using a high-quality frozen 2:4 mask.
| Sparsity | Method | Pattern | LLaMA-2 7B | LLaMA-3.1 8B | ||
| Acc (% ) | PPL () | Acc (% ) | PPL () | |||
| 0% | Dense | - | 54.61 | 5.12 | 60.31 | 5.84 |
| 50% | Magnitude | 2:4 | 43.44 | 54.39 | 35.93 | 765.92 |
| Wanda | 2:4 | 44.30 | 11.15 | 41.77 | 21.29 | |
| SparseGPT | 2:4 | 45.09 | 10.12 | 45.53 | 15.11 | |
| Thanos | 2:4 | 44.80 | 11.19 | 45.72 | 16.09 | |
| ProxSparse | 2:4 | 45.92 | 9.18 | 45.14 | 15.17 | |
| MaskLLM | 2:4 | 48.62 | 6.78 | 52.80 | 8.58 | |
[…]
Overall, across Tables 1 and 2, PATCH consistently surpasses one-shot methods like Wanda, SparseGPT, and magnitude pruning due to its end-to-end training on large corpora. While MaskLLM also trains end-to-end on a large dataset, its fixed 2:4 sparsity ratio limits achievable accuracy and perplexity. In contrast, PATCH overcomes this limitation with flexible dense tile allocation, achieving accuracy gains and perplexity reductions from 45% to 25% sparsity that progressively align with dense model performance. The full per-task accuracy results are provided in Appendix B.
5.2 Understanding the components of PATCH
This subsection examines the design choices driving PATCH’s performance by analyzing its behavior across various configurations on the Qwen-2.5 0.5B model.
Tile size.
We initially assess the impact of tile size on PATCH’s performance, fixing hyperparameters to those optimized for 128128 tiles. Table 4 reveals that tiles maximize model quality through finer sparse-dense control, though larger tile sizes show minimal variation, suggesting robustness. However, smaller tiles may hinder hardware efficiency, requiring a balance with hardware specifications.
128 | 64 | 32 | 16 | 8 | 4 | ||
| 45% | 14.57 | 14.66 | 14.70 | 14.67 | 14.70 | 14.55 | ||
| 35% | 13.84 | 14.08 | 14.15 | 14.03 | 14.01 | 13.72 | ||
| 25% | 13.47 | 13.54 | 13.52 | 13.53 | 13.40 | 13.11 |
Global | Layer-wise | ||
| 45% | 14.57 | 15.17 | ||
| 35% | 13.84 | 14.48 | ||
| 25% | 13.47 | 13.95 |
Joint vs. tile-only mask search.
We then analyze the impact of fixing the 2:4 masks and optimizing only tile masks. Table 5 shows that among frozen 2:4 masks, MaskLLM provides the strongest results. On the other hand, one-shot pruning methods perform comparably at higher sparsity levels but diverge at lower sparsity, with SparseGPT emerging as the best overall. When comparing against our full approach, joint optimization of both tile and 2:4 masks consistently outperforms tile-only training across sparsity ratios. Nevertheless, tile-only training remains a practical alternative for larger models in resource-constrained settings, as also reflected in Table 2.
| Sparsity (0.5B) | MaskLLM |
Wanda | Magnitude | PATCH | ||
| 45% | 15.06 | 21.84 | 21.83 | 21.33 | 14.57 | ||
| 35% | 14.55 | 17.29 | 17.96 | 19.90 | 13.84 | ||
| 25% | 14.17 | 14.89 | 15.09 | 16.05 | 13.47 |
Sparsity allocation.
We analyze how sparsity is allocated across transformer blocks under a global target. Across models, deeper transformer blocks are pruned far less, while the initial blocks also tend to receive lighter pruning depending on the architecture. By contrast, the middle blocks consistently absorb most of the sparsity, suggesting that they contain more redundancy (Figure 2). We compare this flexible allocation to enforcing sparsity uniformly at the layer level. As shown in Table 4, global targets deliver better results by pruning more aggressively in redundant layers while preserving capacity in sensitive ones. In contrast, layer-wise targets impose uniform sparsity that can over-prune critical components (Li et al., 2024b; Xu et al., 2024; Li et al., 2024a; Yin et al., 2025).
On top of variation across depth, sparsity is also distributed unevenly across the individual linear layers within each transformer block. Figure 3 breaks down the allocation into the query, key, value, and output matrices of the attention module, as well as the up, gate, and down matrices of the MLP for the Qwen 2.5 0.5B model. The up, gate, and down layers absorb most of the sparsity and largely explain the overall allocation pattern seen in Figure 2. In contrast, the attention module is treated as more critical. The key and value matrices are never pruned, while the output matrix shows moderate pruning at higher global sparsity targets. The query matrix is pruned the most, suggesting it is the least important within the attention submodule. The distributions for the Gemma-3-1B and LLaMA-3.2-1B models are provided in Appendix E, where the same pattern is observed.
5.3 Combination with other compression methods
LLM compression relies on three orthogonal methods—sparsity, quantization, and low-rank approximation—which can be combined. While this work focuses on sparsity, this section demonstrates how PATCH integrates with these other techniques.
Quantization.
Quantization reduces memory and accelerates computation by lowering numerical precision on hardware optimized for low bitwidths.
Low-rank approximation.
Low-rank methods complement sparsity and quantization by reintroducing a small number of parameters to recover accuracy, with SLiM (Mozaffari et al., 2025a) as a leading one-shot technique.
Table 6 reports results on LLaMA-2 7B and LLaMA-3.1 8B, showing that combining sparsity, quantization, and low-rank approximation enables controllable tradeoffs between compression ratio and model quality, highlighting PATCH ’s versatility.
| Sparsity | Method | Pattern | LoRA | LLaMA-2-7B | LLaMA-3.1-8B | ||
| Acc (% ) | PPL () | Acc (% ) | PPL () | ||||
| 0% | Dense | - | - | 56.44 | 5.12 | 59.85 | 5.84 |
| 50% | MaskLLM | 2:4 | - | 47.98 | 7.64 | 51.12 | 9.92 |
[…]
5.4 Speedup and memory savings
We evaluate the inference efficiency of the LLaMA-2 7B model pruned with PATCH using the STOICC (Rafii et al., 2025) compiler. With a batch size of 16 on an A6000 GPU, we observe end-to-end throughput improvements of 1.18, 1.27, and 1.38 at sparsity levels of 25%, 35%, and 45%, respectively, compared to the dense baseline. At the same sparsity levels, the model’s GPU memory footprint during inference is also reduced, dropping to 0.76×, 0.68×, and 0.59× of the fully dense model, respectively. These results underscore the trade-off between accuracy retention and the computational savings enabled by sparsity.
6 Conclusion
We introduced PATCH, a hybrid sparsity framework that bridges the gap between unstructured and 2:4 sparsity for large language models. By partitioning weight matrices into tiles designated as either dense or 2:4 sparse, PATCH enables adaptive sparsity ratios between 0% and 50%, balancing accuracy and acceleration.
Experiments across models up to 8B parameters show that PATCH consistently improves accuracy over state-of-the-art 2:4 pruning methods while achieving up to 1.38 end-to-end speedup on consumer grade GPUs. These results demonstrate the promise of hybrid sparsity as a practical approach to efficient LLM inference and motivate future work on broader sparsity formats, integration with quantization, and co-design with hardware kernels.
7 Acknowledgments
This work was also supported in part by NSERC Discovery Grants (RGPIN-06516, DGECR00303), the Canada Research Chairs program, Ontario Early Researcher award, the Canada Research Chairs program, the Ontario Early Researcher Award, and the Digital Research Alliance of Canada (www.alliancecan.ca). We extend our gratitude towards Ray Hung for assistance with the results, and Victor Kamel and Arya Rafii for their help in integrating STOICC into our work.
Appendix A STOICC Integration
Triton (Tillet et al., 2019) enables developers to write efficient GPU kernels with a Python-like syntax, but it natively supports only dense matrix operations and cannot handle sparsity. To accelerate the mixed-tile format produced by PATCH, we employ the STOICC compiler (Rafii et al., 2025). STOICC extends Triton with a sparse code-generation backend that allows tiles within a matrix to be either dense or sparse, enabling mixed execution within a single matrix multiplication.
We rely on STOICC’s inspector to autotune both tile sizes and execution schedules (i.e., alternative kernel execution schemes such as split- parallelism) for the prefill and decoding stages of LLM inference. Matrix compression and metadata generation are determined by the chosen tile size, which must remain consistent across both stages. To address this, we first autotune the decoding stage, which is the primary bottleneck of autoregressive generation, since it is executed once per generated token (e.g., 128 times for 128 new tokens), unlike the single pass of prefill. The optimal tile size identified for decoding are then fixed and reused for prefill, where we perform a second round of autotuning over the remaining independent parameters.
In contrast, for fully 2:4 sparse matrices, compression is independent of the block size, so they can be autotuned in the same way as dense kernels in Triton without this coupling constraint.
The pseudocode outlining this process, including the handling of dense, fully 2:4 sparse, and mixed-sparsity modules, is provided in PseudoCode 1.
Table 7 reports the measured throughput (tokens processed per second) of LLaMA-2 7B at sparsity levels of 45%, 35%, and 25% with a batch size of 16 on an A6000 GPU. To reduce CPU overhead from launching Triton kernels in PyTorch, we executed generation through CUDA graphs, capturing both the prefill and decoding stages. With sparsity ratios between 25% and 45%, our heterogeneous approach achieves 1.18–1.38 end-to-end acceleration over the dense baseline. We also report timings on A100 in Table 8.
| Sparsity | Prefill length | Tokens generated | Throughput (tok/s) | Speedup vs. dense |
| 0% | 128 | 128 | 1023.80 | 1.00 |
| 25% | 128 | 128 | 1212.79 | 1.18 |
| 35% | 128 | 128 | 1304.46 | 1.27 |
| 45% | 128 | 128 | 1410.20 | 1.38 |
| 0% | 128 | 1024 | 435.42 | 1.00 |
| 25% | 128 | 1024 | 493.33 | 1.13 |
| 35% | 128 | 1024 | 515.39 | 1.18 |
| 45% | 128 | 1024 | 542.87 | 1.25 |
| Sparsity | Prefill length | Tokens generated | Throughput (tok/s) | Speedup vs. dense |
| 0% | 128 | 128 | 1876.24 | 1.00 |
| 25% | 128 | 128 | 2002.02 | 1.07 |
| 35% | 128 | 128 | 2088.98 | 1.11 |
| 45% | 128 | 128 | 2180.88 | 1.16 |
| 0% | 128 | 1024 | 812.55 | 1.00 |
| 25% | 128 | 1024 | 864.66 | 1.06 |
| 35% | 128 | 1024 | 885.90 | 1.09 |
| 45% | 128 | 1024 | 907.12 | 1.12 |
Appendix B Per task results
This appendix provides detailed per-task accuracy results for the models evaluated in Section 5, covering eight zero-shot downstream tasks: MMLU, PIQA, ARC-Easy, ARC-Challenge, Winogrande, OpenbookQA, RACE, and Hellaswag. The results are presented for each model at various sparsity levels and pruning methods, including our proposed PATCH and PATCH variants, alongside baseline methods such as Magnitude, Wanda, SparseGPT, Thanos, ProxSparse, and MaskLLM. These tables complement the average accuracy and perplexity results reported in Tables 1 and 2 of the main paper, offering a granular view of model performance across individual tasks.
For smaller models (Qwen-2.5 0.5B, LLaMA-3.2 1B, and Gemma-3 1B), we report results using the PATCH variant, which jointly optimizes dense tile locations and sparsity patterns within sparse tiles. For larger models (LLaMA-2 7B and LLaMA-3.1 8B), we report results using the memory-efficient PATCH variant, which optimizes dense tile selections with a fixed 2:4 sparsity mask. The per-task accuracies highlight the effectiveness of our approaches in maintaining robust performance across diverse tasks, even at high sparsity levels, compared to baseline methods.
The following tables detail the per-task accuracies for each model:
Qwen-2.5 0.5B: Table 9 presents the per-task accuracies for the PATCH variant and baselines at 0% and 50% sparsity, with PATCH evaluated at 25%, 35%, and 45% sparsity.
LLaMA-2 7B: Table 10 shows the per-task accuracies for the PATCH variant and baselines, with PATCH evaluated at 25%, 35%, and 45% sparsity.
LLaMA-3.1 8B: Table 11 provides the per-task accuracies for the PATCH variant and baselines, with PATCH at 25%, 35%, and 45% sparsity.
LLaMA-3.2 1B: Table 12 reports the per-task accuracies for the PATCH variant and baselines, with PATCH at 25%, 35%, and 45% sparsity.
Gemma-3 1B: Table 13 details the per-task accuracies for the PATCH variant and baselines, with PATCH at 25%, 35%, and 45% sparsity.
These results enable a deeper analysis of the task-specific performance trends, demonstrating the flexibility and robustness of PATCH and PATCH in achieving high accuracy across diverse tasks while maintaining hardware-friendly sparsity patterns.
| Sparsity | Method | Pattern | MMLU | PIQA | ARC-E | ARC-C | WinoG. | OBQA | RACE | HellaS. | Avg |
| 0% | Dense | - | 47.71 | 70.24 | 64.48 | 29.52 | 56.20 | 24.20 | 35.02 | 40.63 | 46.00 |
| 50% | Magnitude | 2:4 | 23.00 | 54.24 | 31.23 | 19.20 | 49.96 | 13.60 | 23.44 | 26.59 | 30.16 |
| Wanda | 2:4 | 24.43 | 58.71 | 43.18 | 17.75 | 51.62 | 12.20 | 26.32 | 29.58 | 32.97 | |
| SparseGPT | 2:4 | 22.93 | 60.77 | 46.60 | 20.82 | 52.88 | 14.00 | 29.57 | 30.93 | 34.81 | |
| Thanos | 2:4 | 22.97 | 60.17 | 45.37 | 19.20 | 53.59 | 15.20 | 31.00 | 31.31 | 34.85 | |
| ProxSparse | 2:4 | 23.00 | 57.34 | 40.53 | 18.26 | 48.62 | 14.00 | 25.65 | 29.02 | 32.05 | |
| MaskLLM | 2:4 | 25.11 | 67.03 | 56.57 | 23.98 | 52.57 | 20.20 | 33.30 | 35.90 | 39.33 | |
[…]
| Sparsity | Method | Pattern | MMLU | PIQA | ARC-E | ARC-C | WinoG. | OBQA | RACE | HellaS. | Avg |
| 0% | Dense | - | 41.82 | 78.07 | 76.35 | 43.52 | 69.06 | 31.40 | 39.52 | 57.13 | 54.61 |
| 50% | Magnitude | 2:4 | 25.82 | 70.02 | 61.78 | 30.12 | 61.01 | 21.80 | 31.48 | 45.45 | 43.44 |
| Wanda | 2:4 | 25.80 | 71.00 | 63.80 | 30.29 | 61.09 | 25.20 | 35.50 | 41.75 | 44.30 | |
| SparseGPT | 2:4 | 26.17 | 70.73 | 63.80 | 30.63 | 65.04 | 24.00 | 37.13 | 43.18 | 45.09 | |
| Thanos | 2:4 | 25.27 | 70.78 | 63.43 | 30.97 | 64.56 | 23.80 | 36.46 | 43.11 | 44.80 | |
| ProxSparse | 2:4 | 26.77 | 71.60 | 65.70 | 33.02 | 62.90 | 24.20 | 35.31 | 47.84 | 45.92 | |
| MaskLLM | 2:4 | 27.65 | 74.76 | 69.44 | 35.58 | 65.04 | 26.80 | 38.56 | 51.15 | 48.62 | |
[…]
| Sparsity | Method | Pattern | MMLU | PIQA | ARC-E | ARC-C | WinoG. | OBQA | RACE | HellaS. | Avg |
| 0% | Dense | - | 63.57 | 80.09 | 81.44 | 51.37 | 73.48 | 33.40 | 39.14 | 60.02 | 60.31 |
| 50% | Magnitude | 2:4 | 23.06 | 63.82 | 45.33 | 25.94 | 53.91 | 15.20 | 26.70 | 33.49 | 35.93 |
| Wanda | 2:4 | 27.85 | 68.88 | 58.33 | 26.71 | 60.93 | 19.00 | 33.78 | 38.70 | 41.77 | |
| SparseGPT | 2:4 | 31.82 | 70.46 | 63.85 | 31.74 | 64.56 | 21.60 | 37.22 | 42.99 | 45.53 | |
| Thanos | 2:4 | 34.23 | 70.40 | 63.13 | 31.40 | 63.61 | 23.20 | 37.03 | 42.75 | 45.72 | |
| ProxSparse | 2:4 | 29.89 | 71.71 | 62.63 | 33.28 | 58.56 | 23.80 | 35.22 | 46.03 | 45.14 | |
| MaskLLM | 2:4 | 42.47 | 77.04 | 73.15 | 40.19 | 68.43 | 28.80 | 38.28 | 54.04 | 52.80 | |
[…]
| Sparsity | Method | Pattern | MMLU | PIQA | ARC-E | ARC-C | WinoG. | OBQA | RACE | HellaS. | Avg |
| 0% | Dense | - | 37.57 | 74.54 | 65.53 | 31.32 | 60.62 | 26.40 | 37.89 | 47.76 | 47.70 |
| 50% | Magnitude | 2:4 | 23.31 | 53.81 | 27.74 | 18.94 | 51.38 | 11.80 | 24.02 | 26.26 | 29.66 |
| Wanda | 2:4 | 22.90 | 58.11 | 37.08 | 19.20 | 49.09 | 13.20 | 25.17 | 28.11 | 31.61 | |
| SparseGPT | 2:4 | 22.93 | 61.43 | 45.03 | 22.35 | 54.93 | 15.80 | 29.86 | 32.08 | 35.55 | |
| Thanos | 2:4 | 23.12 | 62.40 | 44.91 | 21.76 | 54.30 | 16.00 | 31.10 | 32.09 | 35.71 | |
| ProxSparse | 2:4 | 22.96 | 60.83 | 39.44 | 20.31 | 51.54 | 16.80 | 25.17 | 31.37 | 33.55 | |
| MaskLLM | 2:4 | 26.28 | 69.10 | 57.41 | 25.85 | 55.48 | 21.40 | 32.82 | 39.94 | 41.04 | |
[…]
| Sparsity | Method | Pattern | MMLU | PIQA | ARC-E | ARC-C | WinoG. | OBQA | RACE | HellaS. | Avg |
| 0% | Dense | - | 24.95 | 75.03 | 71.84 | 34.90 | 58.64 | 28.60 | 34.83 | 47.26 | 47.01 |
| 50% | Magnitude | 2:4 | 23.08 | 59.79 | 37.29 | 17.66 | 50.59 | 14.00 | 22.87 | 27.97 | 31.66 |
| Wanda | 2:4 | 23.96 | 59.52 | 48.02 | 18.34 | 51.22 | 14.20 | 27.85 | 30.18 | 34.16 | |
| SparseGPT | 2:4 | 23.62 | 62.79 | 49.83 | 19.03 | 51.54 | 15.20 | 30.62 | 31.99 | 35.58 | |
| Thanos | 2:4 | 23.44 | 62.24 | 48.86 | 18.34 | 50.12 | 15.60 | 30.81 | 31.28 | 35.09 | |
| ProxSparse | 2:4 | 23.10 | 64.25 | 50.72 | 21.59 | 53.43 | 18.00 | 29.09 | 32.86 | 36.63 | |
| MaskLLM | 2:4 | 25.03 | 69.91 | 60.27 | 27.65 | 56.27 | 21.20 | 34.55 | 39.84 | 41.84 | |
[…]
Appendix C Tile Transfer Learning
We also test whether initializing tile logits with priors from one-shot pruning methods improves performance, as done in MaskLLM (Fang et al., 2024). In our case, the initialization is derived from one-shot pruning with unstructured sparsity. We initialize tiles that retain more nonzeros after unstructured pruning with positive logits (favoring dense assignment), while the remaining tiles receive negative logits, controlled by a strength parameter. The number of tiles initialized as dense is selected such that the overall layer-wise sparsity target is satisfied. As shown in Table 14, the choice of prior has little impact on final performance: all priors yield nearly identical perplexity, with random initialization often performing best. This is likely because the global sparsity target enables dynamic reallocation of sparsity across layers during training, overriding the effect of any fixed initialization. For consistency with prior work, we adopt SparseGPT initialization in all experiments.
| Sparsity (0.5B) | Nothing | SparseGPT | Wanda | Magnitude | Random |
| 45% | 14.80 | 14.57 | 14.50 | 14.48 | 14.51 |
| 35% | 13.97 | 13.84 | 13.87 | 13.85 | 13.79 |
| 25% | 13.47 | 13.47 | 13.37 | 13.44 | 13.33 |
Appendix D Implementation details and hyperparameters
We train all masks using the HuggingFace Trainer API (Wolf et al., 2020) for 2000 steps with a global batch size of 256 and a sequence length of 4096, resulting in 2B tokens processed from the SlimPajama corpus (Soboleva et al., 2023).
Training is accelerated via data parallelism across a single node with 4 H100 GPUs. In this setup, PATCH requires 4.5 and 6 GPU hours on the 0.5B and 1B models, respectively, while PATCH requires 21 and 24 GPU hours on the 7B and 8B models.
The hyperparameters for PATCH and PATCH are summarized in Table 15, tuned on Qwen-2.5-0.5B. For the 2:4 mask parameters, we follow the configuration from MaskLLM (Fang et al., 2024).
| Sparsity | Method | Optimizer | Logits Init | Gumbel Scaling | Gumbel | Prior(Strength) | Sparse Reg. | Weight Reg. |
| 25% | PATCH | Adam(0.001) | SparseGPT() | 7 | 10 | |||
| 35% | PATCH | Adam(0.001) | SparseGPT() | 7 | 10 | |||
| 45% | PATCH | Adam(0.001) | SparseGPT() | 7 | 10 | |||
| 25% | PATCH | Adam(0.0001) | SparseGPT() | 3 | 0.1 | |||
| 35% | PATCH | Adam(0.0001) | SparseGPT() | 3 | 0.1 | |||
| 45% | PATCH | Adam(0.0001) | SparseGPT() | 3 | 0.1 |
Appendix E Additional layer-wise sparsity distributions
In this appendix, we provide the sparsity distributions for the Gemma-3-1B (Figure 4) and Llama-3.2-1B (Figure 5) models, as referenced in the main text. Similar to the Qwen-2.5 0.5B model, the patterns observed here indicate that MLP layers (up, gate, and down matrices) are pruned more aggressively, absorbing the majority of sparsity. In contrast, the self-attention layers are treated as more critical, with key and value matrices remaining largely dense or unpruned, while the query matrix experiences the highest pruning within the attention submodule, and the output matrix shows moderate pruning under higher global sparsity targets. This consistent behavior across models underscores the redundancy in MLP components and the sensitivity of attention mechanisms.
Appendix F Related work
F.1 Pruning methods
Pruning is one of the most widely studied approaches for compressing deep neural networks, with the goal of removing redundant parameters while preserving accuracy. Classical pruning methods can be broadly categorized into local (layer-wise) and global (end-to-end) strategies.
Local pruning.
Local approaches prune each layer independently, typically by minimizing reconstruction error within that layer. A seminal example is Optimal Brain Surgeon (OBS) (Hassibi et al., 1993; Frantar & Alistarh, 2022), which leverages second-order information to identify and remove weights while updating the remaining parameters to compensate for loss. While highly principled, the quadratic cost of computing and inverting the Hessian makes OBS infeasible for large models.
Recent work adapts these ideas to LLM-scale pruning. SparseGPT (Frantar & Alistarh, 2023) formulates layer-wise pruning as a sparse regression problem, enabling efficient approximations of OBS that scale to billion-parameter models. Thanos (Ilin & Richtarik, 2025) further improves accuracy by employing multi-column approximations to reduce error accumulation. Wanda (Sun et al., 2023), on the other hand, discards explicit weight updates and instead uses a simple magnitude-activation criterion with calibration data, yielding competitive quality with extremely fast runtimes. Despite their efficiency, local methods often suffer from limited capacity to recover accuracy since pruning decisions ignore cross-layer dependencies.
Global pruning.
Global approaches aim to jointly optimize pruning decisions across layers, typically leading to better overall trade-offs. Optimal Brain Damage (OBD) (LeCun et al., 1989) is an early global method that estimates weight saliency using the diagonal Hessian. Extensions such as WoodFisher (Singh & Alistarh, 2020) approximate the Hessian via Kronecker factorizations, making computation more tractable but still challenging for modern LLMs (Mozaffari et al., 2023).
More recent approaches bypass costly second-order computations. MaskLLM (Fang et al., 2024) formulates pruning as a binary classification task (keep vs. prune) and solves it using standard optimizers such as AdamW (Loshchilov, 2017), achieving strong results even under hardware-friendly structured sparsity (e.g., ). ProxSparse (Liu et al., 2025) instead adopts a proximal regularization framework, reducing the overhead of MaskLLM while trading off some pruning accuracy. These works highlight the tension between pruning quality and efficiency: global methods often achieve higher accuracy but remain more computationally expensive than simple one-shot local pruning.
F.2 Complementary compression techniques
Beyond pruning, several orthogonal compression techniques are widely used and can be combined with sparsity for additional gains. Quantization reduces the bit precision of parameters and activations, e.g., from 32-bit floating point to 8- or 4-bit integers, thereby reducing memory footprint and accelerating inference (Gholami et al., 2022; Rokh et al., 2023).
Low-rank adaptation methods decompose weight matrices into smaller factors, effectively reducing parameter counts while maintaining expressivity. Recent approaches such as LQ-LoRA (Guo et al., 2023), SLiM (Mozaffari et al., 2025a), and SLoPe (Mozaffari et al., 2025b) demonstrate that low-rank structures can be used both for efficient fine-tuning and for direct model compression.
Finally, knowledge distillation (Gou et al., 2021) transfers knowledge from a large teacher model to a smaller student, yielding compact models that retain much of the teacher’s performance. These methods are complementary to pruning, and hybrid frameworks that integrate sparsity, quantization, and low-rank factorization represent a promising direction for achieving high compression ratios without sacrificing accuracy.
Appendix G Comparison with unstructured sparsity
In this section, we compare the quality of the models pruned with PATCH against other unstructured sparsity methods. Table 16 summarizes the average accuracy of the models across eight downstream tasks and the model perplexity on WikiText2 dataset. The results indicate that while unstructured sparsity consistently outperforms the hybrid sparsity, the gap between the two is not significant, showing that PATCH is helping bridging the gap between unstructured sparsity and semi-structured sparsity.
| Sparsity | Method | Pattern | Qwen-2.5 0.5B | LLaMA-3.2 1B | Gemma-3 1B | LLaMA-2 7B | LLaMA-3.1 8B | |||||
| Acc (% ) | PPL () | Acc (% ) | PPL () | Acc (% ) | PPL () | Acc (% ) | PPL () | Acc (% ) | PPL () | |||
[…]
| 45% | Wanda | Unstructured | 41.45 | 18.81 | 40.76 | 16.56 | 42.87 | 25.38 | 52.72 | 6.36 | 55.67 | 8.24 |
| 45% | SparseGPT | Unstructured | 42.31 | 17.65 | 42.66 | 15.01 | 43.52 | 22.26 | 52.77 | 6.46 | 56.70 | 8.21 |
[…]
| 35% | Wanda | Unstructured | 43.46 | 15.04 | 44.60 | 11.95 | 45.50 | 16.98 | 54.37 | 5.87 | 58.68 | 7.02 |
| 35% | SparseGPT | Unstructured | 44.66 | 14.79 | 45.62 | 11.68 | 45.45 | 16.92 | 54.18 | 5.92 | 58.81 | 7.07 |
[…]
| 25% | Wanda | Unstructured | 45.70 | 13.70 | 46.50 | 10.46 | 46.56 | 15.14 | 54.60 | 5.65 | 59.80 | 6.54 |
| 25% | SparseGPT | Unstructured | 45.28 | 13.63 | 46.52 | 10.42 | 46.37 | 15.05 | 54.71 | 5.68 | 59.52 | 6.55 |
Appendix H Language model usage in paper
We used language models to enhance the readability of the manuscript, correct grammatical and typographical errors, and ensure conformity with the ICLR author guidelines. Beyond their application in benchmark evaluations and experimental procedures, language models were not employed in any other aspect of this study.
Appendix I Reproducibility statement
To support reproducibility, we release a repository linked in the abstract footnote that contains our implementation, training scripts, and evaluation pipeline. The paper outlines the method in 3 and provides a thorough experimental description in 5. Appendix D discusses the hyperparameter values in our work and additional information about our implementation. These materials collectively allow others to replicate our experiments and validate the claims made in the paper.